OpenStudy (anonymous):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4^-x and y = 2^(x + 3) intersect are the solutions of the equation 4^-x = 2^(x + 3). Part B: Make tables to find the solution to 4^-x = 2^(x + 3). Take the integer values of x between -3 and 3. Part C: How can you solve the equation 4^-x = 2^(x + 3) graphically?

OpenStudy (anonymous):

for part A, the points on the graph of both are \((x,y)\) and in once case it is \((x, 4^{-x})\) and in the other it is \((x,2^{x+3})\) where they intersect both the \(x\) and the \(y\) values will be the same, so \(4^{-x}\) will be equal to \(2^{x+3}\) at that point

OpenStudy (anonymous):

Thanks :). Can you help me with part 2? I don't know how to make the tables.

OpenStudy (anonymous):

it means pick \(x\) as the first number and compute \(y\) for the second for example, if \(x=-3\) then \(y=4^{-x}=4^{-(-3)}=4^3=64\) and \[2^{x+3}=2^{2+3}=2^5=32\]

OpenStudy (anonymous):

actually i made a mistake there if \(x=-3\) then \[2^{x+2}=2^{-3+2}=2^{-1}=\frac{1}{2}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

damn another mistake, it is \(2^{x+3}\) not \(2^{x+2}\) scratch that \[2^{-3-3}=2^0=1\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Wait, so would the tables look like this?: Table for y = 4^-x: x y -3 64 -2 16 -1 4 0 1 1 0.25 2 0.0625 3 0.015625 Table for y = 2^(x+3): x y -3 1 -2 2 -1 3 0 8 1 16 2 32 3 64

OpenStudy (anonymous):

@satellite73 ?

OpenStudy (anonymous):

Wait sorry like this: Table for y = 4^-x: x y -3 64 -2 16 -1 4 0 1 1 0.25 2 0.0625 3 0.015625 Table for y = 2^(x+3): x y -3 1 -2 2 -1 4 0 8 1 16 2 32 3 64

OpenStudy (anonymous):

@wio @shamim ? are you guys there? could you please help me?